OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10 i will friend and medal

OpenStudy (alexandervonhumboldt2):

g(x)=20 g(x)=2*(x-4) => 20=2*(x-4) can you solve the rest?

OpenStudy (anonymous):

2x-8

OpenStudy (alexandervonhumboldt2):

yes so we have 2x-8=20 2x=28 can you solve for x?

OpenStudy (calculusxy):

okay so you need to find the value of x by \(20 = 2(x - 4)\) \(20 = 2x - 8\) \(28 = 2x\) solve for x, just like how @AlexandervonHumboldt2, said.

OpenStudy (calculusxy):

\[\large \frac{ 28 }{ 2 } = \frac{ 2x }{ 2 }\]

OpenStudy (anonymous):

14??

OpenStudy (calculusxy):

correct!

OpenStudy (anonymous):

thank you!

OpenStudy (calculusxy):

ur welcome :)